# Graph Retrieval-Augmented Generation: A Survey

### v1

@@ 9.2.2. Biomedical
Recently, GraphRAG techniques are increasingly applied in biomedical question answering systems, achieving advanced medical decision-making performance. In this area, each disease is associated with specific symptoms, and every medication contains certain active ingredients that target and treat particular diseases. Some researchers (Li et al., 2024e; Delile et al., 2024) construct KGs for specific task scenarios, while others (Jiang et al., 2024a; Yang et al., 2024a; Wen et al., 2024) utilize open-source knowledge graphs such as CMeKG and CPubMed-KG as retrieval sources. Existing methods generally begin with non-parametric retrievers for initial search, followed by designing methods to filter retrieved content through reranking (Jiang et al., 2024a; Yang et al., 2024a; Wen et al., 2024; Li et al., 2024e; Delile et al., 2024). Additionally, some approaches propose rewriting model inputs using retrieved information to enhance generation effectiveness (Li et al., 2024e).
@@ 9.2.5. Legal
In legal contexts, extensive citation connections exist between cases and judicial opinions, as judges frequently reference previous opinions when making new decisions. This naturally creates a structured graph where nodes represent opinions, opinion clusters, dockets, and courts, and edges encompass relationships such as “opinion-citation”, “opinion-cluster”, “cluster-docket”, and “docket-court”. The application of GraphRAG in legal scenario could aid lawyers and legal researchers in various tasks such as case analysis and legal consultation.
@@ 9.2.6. Others
In addition to the above applications, GraphRAG is also applied to other real-world scenarios such as intelligence report generation (Ranade and Joshi, 2023) and patent phrase similarity detection (Peng and Yang, 2024)Ranade and Joshi (2023) first construct an Event Plot Graph (EPG) and retrieve the critical aspects of the events to aid the generation of intelligence reports. Peng and Yang (2024) create a patent-phrase graph and retrieve the ego-network of the given patent phrase to assist the judgment of phrase similarity.
@@ 9.3.1. Benchmarks
The benchmarks used to evaluate the performance of the GraphRAG system can be divided into two categories. The first category is the corresponding datasets of downstream tasks. We summarize the benchmarks and papers tested with them according to the classification in Section 9.1, details of which are shown in Table 1. The second category consists of benchmarks specifically designed for the GraphRAG systems. These benchmarks usually cover multiple task domains to provide a comprehensive test result. For example, STARK (Wu et al., 2024b) benchmarks LLM Retrieval on semi-structured knowledge bases covering three domains, including product search, academic paper search, and queries in precision medicine to access the capacity of current GraphRAG systems. He et al. (2024) propose a flexible question-answering benchmark targeting real-world textual graphs, named GraphQA, which is applicable to multiple applications including scene graph understanding, commonsense reasoning, and knowledge graph reasoning. Graph Reasoning Benchmark (GRBENCH) (Jin et al., 2024b) is constructed to facilitate the research of augmenting LLMs with graphs, which contains 1,740 questions that can be answered with the knowledge from 10 domain graphs. CRAG (Yang et al., 2024b) provides a structured query dataset, with additional mock APIs to access information from underlying mock KGs to achieve fair comparison.
@@ 10.1. Dynamic and Adaptive Graphs
Most GraphRAG methods (Edge et al., 2024; Fu et al., 2020; Lan et al., 2023, 2021; Yani and Krisnadhi, 2021; Luo et al., 2024a) are built upon static databases; however, as time progresses, new entities and relationships inevitably emerge. Rapidly updating these changes is both promising and challenging. Incorporating updated information is crucial for achieving better results and addressing emerging trends that require current data. Developing efficient methods for dynamic updates and real-time integration of new data will significantly enhance the effectiveness and relevance of GraphRAG systems.

### v2

@@ 9.2.2. Biomedical
Recently, GraphRAG techniques are increasingly applied in biomedical question answering systems, achieving advanced medical decision-making performance. In this area, each disease is associated with specific symptoms, and every medication contains certain active ingredients that target and treat particular diseases. Some researchers (Li et al., 2024e; Delile et al., 2024; Wu et al., 2024d) construct KGs for specific task scenarios, while others (Jiang et al., 2024a; Yang et al., 2024a; Wen et al., 2024) utilize open-source knowledge graphs such as CMeKG and CPubMed-KG as retrieval sources. Existing methods generally begin with non-parametric retrievers for initial search, followed by designing methods to filter retrieved content through re-ranking (Jiang et al., 2024a; Yang et al., 2024a; Wen et al., 2024; Li et al., 2024e; Delile et al., 2024). Additionally, some approaches propose rewriting model inputs using retrieved information to enhance generation effectiveness (Li et al., 2024e).
@@ 9.2.5. Legal
In legal contexts, extensive citation connections exist between cases and judicial opinions, as judges frequently reference previous opinions when making new decisions. This naturally creates a structured graph where nodes represent opinions, opinion clusters, dockets, and courts, and edges encompass relationships such as “opinion-citation”, “opinion-cluster”, “cluster-docket”, and “docket-court”. The application of GraphRAG in legal scenarios could aid lawyers and legal researchers in various tasks such as case analysis and legal consultation.
@@ 9.2.6. Others
In addition to the above applications, GraphRAG is also applied to other real-world scenarios such as intelligence report generation (Ranade and Joshi, 2023), patent phrase similarity detection (Peng and Yang, 2024) and software understanding (Alhanahnah et al., 2024)Ranade and Joshi (2023) first construct an Event Plot Graph (EPG) and retrieve the critical aspects of the events to aid the generation of intelligence reports.  Peng and Yang (2024) create a patent-phrase graph and retrieve the ego network of the given patent phrase to assist the judgment of phrase similarity.   Alhanahnah et al. (2024) propose a Chatbot to understand properties about dependencies in a given software package, which first automatically constructs the dependency graph and then the user can ask questions about the dependencies in the dependency graph.
@@ 9.3.1. Benchmarks
The benchmarks used to evaluate the performance of the GraphRAG system can be divided into two categories. The first category is the corresponding datasets of downstream tasks. We summarize the benchmarks and papers tested with them according to the classification in Section 9.1, details of which are shown in Table 1. The second category consists of benchmarks specifically designed for the GraphRAG systems. These benchmarks usually cover multiple task domains to provide a comprehensive test result. For example, STARK (Wu et al., 2024c) benchmarks LLM Retrieval on semi-structured knowledge bases covering three domains, including product search, academic paper search, and queries in precision medicine to access the capacity of current GraphRAG systems.  He et al. (2024) propose a flexible question-answering benchmark targeting real-world textual graphs, named GraphQA, which is applicable to multiple applications including scene graph understanding, commonsense reasoning, and knowledge graph reasoning. Graph Reasoning Benchmark (GRBENCH) (Jin et al., 2024b) is constructed to facilitate the research of augmenting LLMs with graphs, which contains 1,740 questions that can be answered with the knowledge from 10 domain graphs. CRAG (Yang et al., 2024b) provides a structured query dataset, with additional mock APIs to access information from underlying mock KGs to achieve fair comparison.
@@ 10.1. Dynamic and Adaptive Graphs
Most GraphRAG methods (Edge et al., 2024; Fu et al., 2020; Lan et al., 2023, 2021; Yani and Krisnadhi, 2021; Luo et al., 2024a) are built upon static databases; however, as time progresses, new entities and relationships inevitably emerge (Wu et al., 2024a; Cheng et al., 2024; Gao et al., 2024a). Rapidly updating these changes is both promising and challenging. Incorporating updated information is crucial for achieving better results and addressing emerging trends that require current data. Developing efficient methods for dynamic updates and real-time integration of new data will significantly enhance the effectiveness and relevance of GraphRAG systems.